# DualAnoDiff: Dual-Interrelated Diffusion Model for Few-Shot Anomaly Image Generation

### v1

@@ Introduction
$K=W^{(i)}_{Q}\cdot$ $\varphi_{i}(z)\quad V=W^{(i)}_{Q}\cdot\varphi_{i}(z)$ (5)
@@
$+\mathbb{E}_{\mathcal{E}(I),\epsilon^{\ast},t}\left[\left\|\epsilon^{\ast}-\epsilon_{\theta}^{\ast}\left(z^{\prime}_{t},t,\tau_{\theta}(p^{\prime})\right)\right\|_{2}^{2}\right]$ (6)
@@
Table 2: Comparison on pixel-level and image-level anomaly localization on MVTec dataset by training an U-Net on the generated data from DRAEM, PRN, DFMGAN, AnomalyDiffusion and our model.
Experiments
Experiment Settings
Dataset. we conduct experiments on MVTec (Bergmann et al. 2019). We refer to the experimental setup of AnomalyDiffusion (Hu et al. 2024) using 1/3 of the data as the training set, reserving the remaining two-thirds for testing. Implementation details. We train a model for each anomaly type separately and generate 500 anomaly image-mask pairs for the downstream anomaly inspection tasks. More details of the experiment can be found in the supplementary materials. Metric. 1) For generation, we use Inception Score (IS) and Intra-cluster pairwise LPIPS distance (IC-LPIPS) (Ojha et al. 2021) as AnomalyDiffusion (Hu et al. 2024) to measure the generation quality and generation diversity. 2) For anomaly inspection. We use AUROC, Average Precision (AP), and the F1-max score to evaluate the accuracy of anomaly detection and localization.
Comparison in Anomaly Generation
Baseline. 1) We select DRAEM(Zavrtanik, Kristan, and Skočaj 2021), PRN (Zhang et al. 2023), DFMGAN (Duan et al. 2023) and AnomalyDiffusion (Hu et al. 2024) which can generate anomaly image-mask pairs as comparative benchmarks to compare anomaly detection and localization. 2) We choose CDC (Ojha et al. 2021), Crop-Paste, SDGAN (Niu et al. 2020), DefectGAN (Zhang et al. 2021), DFMGAN (Duan et al. 2023) and AnomalyDiffusion (Hu et al. 2024) that can only generate specific anomaly types to compare anomaly generation quality and classification.
@@
zipper 22.76 26.42 27.64 69.51 69.51
Average 39.31 40.55 49.61 66.09 72.26
Table 3: Comparison on anomaly classification accuracy trained on the generated data by the anomaly generation models with a ResNet-18.
Anomaly Generation Quality. Tab. 1 shows the results of the quantification metrics for image quality. For each anomaly category, we allocate one-third of the anomaly data for training and generate 500 anomaly images to compute IS and IC-LPIPS. It demonstrates that our model generates anomaly data with both the highest quality and diversity.
@@
Table 4: Comparison on pixel-level anomaly localization (AUROC/AP) between the simple U-Net trained on our generated dataset and the existing anomaly detection methods with their official codes or pre-trained models.
Moreover, we present the anomaly images generated by several prominent anomaly image generation models in Fig. 5, among which, AnomalyDiffusion is by far the stat-of-art method. It is observed that the images generated by Defect-GAN often do not have anomalies, DFMGAN tends to introduce noise in the image, and the anomaly image-mask pairs lack proper alignment. The state-of-the-art model AnomalyDiffusion may generates an unreasonable mask that is outside the object, and the generated anomaly does not fit the original image enough to make the anomaly look unreal. In contrast, our proposed model demonstrates the ability to generate highly realistic and valid anomalies, even excelling in handling smaller-scale anomalies such as grid-bent. More results are presented in the supplementary material.
Anomaly Generation for Anomaly Detection and Localization. we evaluate the effectiveness of our approach by comparing it with existing methods for anomaly generation in downstream anomaly detection and localization. We calculate pixel-level dimensions and image-level AUROC, AP, and F1-max for each category. The results are presented in Tab. 2, and because the image-level results of AUROC, AP, and F1-max are relatively similar, as well as the limitation of the table width, we only show the results of the most representative AP, and the other results of AUROC and F1-max are shown in the supplementary material. It can be seen that the U-net trained in our generated data reaches the highest AP of 84.5% and F1-max of 78.8%, surpassing the second-ranked AnomalyDiffusion by a margin of 3.1%(AP).
Anomaly Generation for Anomaly Classification. To further assess the quality of the anomaly images generated by our model, we utilize them to train a downstream anomaly classification model. Following the experimental setup of DFMGAN (Duan et al. 2023), we employ ResNet-34 to train on the generated dataset. We then evaluate the classification accuracy on our test dataset as shown in Tab. 3. As evident from the data, our model demonstrates significantly higher accuracy across the majority of categories compared to other models, with an average accuracy improvement of 6.17%, indicating the anomaly data we generate is more realistic.
@@
Comparison with Anomaly Detection Models
To further validate the effectiveness of our model, we compare it with state-of-the-art anomaly detection methods DRAEM(Zavrtanik, Kristan, and Skočaj 2021), SSPCAB (Ristea et al. 2022), CFA (Lee, Lee, and Song 2022), RD4AD (Deng and Li 2022), PatchCore (Roth et al. 2022), MuSc (Li et al. 2024), DevNet (Pang et al. 2021), DRA (Ding, Pang, and Shen 2022) and PRN (Zhang et al. 2023). We utilize their official codes or pre-trained models and evaluate them on the same testing dataset (2/3 anomaly data in the test of Mvtec). As there is no available open-source code for PRN, we rely on the data provided in its research paper. The methods are evaluated based on their pixel-level AUROC and AP scores, as demonstrated in Tab. 4.
@@
Table 5: Comparison on pixel-level anomaly localization on part of categories in MVTec between with/without the Background Compensation Module. Average is the mean of all categories in Mvtec.
Ablation Study
@@
Fig. 6 presents the results of using different prompts in our model. It is evident that using the prompt “a vfx with sks” yields superior performance in terms of object shape, color, and other aspects compared to using the category name directly. More quantitative indicators can be found in the supplementary material. Furthermore, we assess the effectiveness of the Background Compensation Module (BCM), which exhibits more noticeable improvements in the categories of bottle, pill, and toothbrush. The visualization results are depicted in Fig. 4, and the pixel-level localization results are presented in Tab. 5.
Conclusion
@@
In this paper, we present a novel approach, DualAnoDiff, for generating anomalous image-mask pairs. Our method employs a parallel dual-diffusion to simultaneously generate the anomaly image and the corresponding anomaly part image. This ensures a high level of alignment between the generated anomaly image-mask pair and the authenticity of the anomaly. Additionally, to address challenging cases, we introduce a background compensation module that efficiently enhances the model’s fitting capability. Extensive experiments demonstrate its superior performance compared to existing anomaly generation methods. The anomaly data generated by our model effectively enhances the performance of downstream anomaly inspection tasks.

### v3

@@ 3.3 Background Compensation Module
$K=W^{(i)}_{K}\cdot$ $\varphi_{i}(z)\quad V=W^{(i)}_{V}\cdot\varphi_{i}(z)$ (5)
@@
$+\mathbb{E}_{\mathcal{E}(I_{a}),\epsilon^{\ast},t}\left[\left\|\epsilon^{\ast}-\epsilon_{\theta}^{\ast}\left(z^{\prime}_{t},t,\tau_{\theta}(p^{\prime})\right)\right\|_{2}^{2}\right]$ (6)
bottle 1.52 0.04 1.43 0.04 1.57 0.06 1.39 0.07 1.62 0.12 1.58 0.19 2.17 0.43
cable 1.97 0.19 1.74 0.25 1.89 0.19 1.70 0.22 1.96 0.25 2.13 0.41 2.15 0.43
capsule 1.37 0.06 1.23 0.05 1.49 0.03 1.59 0.04 1.59 0.11 1.59 0.21 1.62 0.32
grid 1.97 0.07 2.00 0.12 1.95 0.10 2.01 0.12 1.97 0.13 2.04 0.44 2.13 0.42
hazel_nut 1.97 0.05 1.74 0.21 1.85 0.16 1.87 0.19 1.93 0.24 2.13 0.31 1.94 0.35
leather 1.80 0.07 1.47 0.14 2.04 0.12 2.12 0.14 2.06 0.17 1.94 0.41 1.91 0.35
metal_nut 1.55 0.04 1.56 0.15 1.45 0.28 1.47 0.30 1.49 0.32 1.96 0.30 1.57 0.32
pill 1.56 0.06 1.49 0.11 1.61 0.07 1.61 0.10 1.63 0.16 1.61 0.26 1.82 0.38
screw 1.13 0.11 1.12 0.16 1.17 0.10 1.19 0.12 1.12 0.14 1.28 0.30 1.43 0.36
tile 2.10 0.12 1.83 0.20 2.53 0.21 2.35 0.22 2.39 0.22 2.54 0.55 2.40 0.50
transistor 1.61 0.13 1.39 0.15 1.76 0.13 1.47 0.13 1.64 0.25 1.57 0.34 1.71 0.33
wood 2.05 0.03 1.95 0.23 2.12 0.25 2.19 0.29 2.12 0.35 2.33 0.37 2.24 0.40
zipper 1.30 0.05 1.23 0.11 1.25 0.10 1.25 0.10 1.29 $\underline{0.27}$ 1.39 0.25 2.14 0.37
Average 1.65 0.07 1.51 0.14 1.71 0.13 1.69 0.15 1.72 0.20 1.80 0.32 1.93 0.38
From another perspective, the dual-branch structure is equivalent to explicitly separating the two concepts of the object body and the anomaly in image. However, in the case of some images with backgrounds, two attributes are not enough. model may mix the object body with the solid-color background, thus resulting in the phenomenon shown in Fig. 4(a). The BCM module further decouples the object body from the background by incorporating background information during training, thereby achieving a more stable generation effect in Fig. 4(b).
4 Experiments
4.1 Experiment Settings
Dataset. We conduct experiments on MVTec [1] and follow the experimental setup of AnomalyDiffusion [14] using 1/3 of data for training, reserving the remaining 2/3 for testing.
Implementation Details. We train a model for each anomaly type separately and generate 1000 anomaly image-mask pairs for the downstream anomaly inspection tasks. More details of the experiment can be found in the supplementary materials.
Metric. 1) For generation, we use Inception Score (IS) and Intra-cluster pairwise LPIPS distance (IC-LPIPS) [27] as AnomalyDiffusion [14] to measure the generation quality and generation diversity. 2) For anomaly inspection. We use AUROC, Average Precision (AP), and the F1-max score to evaluate the accuracy of anomaly detection and localization, use Accuracy to evaluate the anomaly classification.
4.2 Comparison in Anomaly Generation
Baseline. 1) We choose CDC [27], Crop-Paste[23], SDGAN [26], DefectGAN [47], DFMGAN [6] and AnomalyDiffusion [14] that can generate specific anomaly types to compare anomaly generation quality and classification. 2) We select DRAEM[46], PRN [48], DFMGAN [6] and AnomalyDiffusion [14] which can generate anomaly image-mask pairs as comparative benchmarks to compare anomaly detection and localization.
@@ 4.2 Comparison in Anomaly Generation
Table 2: Quantitative result for anomaly detection. Comparison on pixel-level anomaly localization and image-level anomaly detection on MVTec dataset by training an U-Net on the generated data from DRAEM, PRN, DFMGAN, AnomalyDiffusion and our model.
@@
zipper 22.76 26.42 27.64 69.51 75.61
Average 39.31 40.55 49.61 66.09 79.67
Table 3: Quantitative result for anomaly classification. Comparison on the classification accuracy(%) trained on the generated data by the anomaly generation models with a ResNet-18. The higher classification accuracy indicates that the generated data is more consistent with the distribution of real data.
Anomaly Generation Quality. Tab. 1 shows the results of the quantification metrics for image quality. For each anomaly category, we allocate 1/3 of the anomaly data for training and generate 1000 anomaly images to compute IS and IC-LPIPS. Through this process, it is clearly demonstrated that our method achieves the best results in both the IS and IC-LPIPS metrics. Moreover, it further shows that our model generates anomaly data with both the highest quality and diversity.
@@
Table 4: Quantitative result for pixel-level anomaly localization (AUROC/AP). We employ a simple U-Net trained on our generated dataset and the existing anomaly detection methods with their official codes or pre-trained models.
Moreover, we present the anomaly images generated by several prominent anomaly image generation models in Fig. 5. It is observed that the images generated by Defect-GAN often do not have anomalies, DFMGAN tends to introduce noise in the image (bottle, capsule), and the anomaly image-mask pairs lack proper alignment (grid). The state-of-the-art model AnomalyDiffusion sometimes generates unreasonable masks outside the objects (grid, screw), and the generated anomaly does not fit the original image enough to make the anomaly look unreal (toothbrush). In contrast, our proposed model demonstrates the ability to generate highly realistic and valid anomalies, even excelling in handling smaller-scale anomalies such as grid-bent. More results are presented in the supplementary material.
Anomaly Generation for Anomaly Detection and Localization. We evaluate the effectiveness of our approach by comparing it with existing methods for anomaly generation in downstream anomaly detection and localization. We calculate pixel-level dimensions and image-level AUROC, AP, and F1-max for each category. The results are presented in Tab. 2, and because the image-level results of AUROC, AP, and F1-max are relatively similar, as well as the space limitation, we only show the results of the most representative AP, and the other results of AUROC and F1-max are shown in the supplementary material. It can be seen that the U-net trained in our generated data reaches the highest AP of 84.5% and F1-max of 78.8%, surpassing the second-ranked AnomalyDiffusion by a margin of 3.1%(AP).
Anomaly Generation for Anomaly Classification. To further assess the quality of the anomaly images generated by our model, we utilize them to train a downstream anomaly classification model. Following the experimental setup of DFMGAN [6], we employ ResNet-18 to train on the generated dataset. We then evaluate the classification accuracy on our test dataset as shown in Tab. 3. As evident from the data, our model demonstrates significantly higher accuracy across the majority of categories compared to other models, with an average accuracy improvement of 13.58%, indicating the anomaly data we generate is more realistic.
@@
4.3 Comparison with Anomaly Detection Models
To further validate the effectiveness of our model, we compare it with state-of-the-art anomaly detection methods DRAEM[46], SSPCAB [32], CFA [19], RD4AD [4], PatchCore [36], MuSc [22], DevNet [28], DRA [5] and PRN [48]. We utilize their official codes or pre-trained models and evaluate them on the same testing dataset (2/3 anomaly data in the test of MVTec). As there is no available open-source code for PRN, we rely on the results provided in its research paper. The methods are evaluated based on their pixel-level AUROC and AP scores, as demonstrated in Tab. 4. It is clearly discernible that upon using the anomaly data generated by our method, even if only the simplest U-Net model is utilized, the result shows a significant increase of 5.9% over that of the traditional method.
4.4 Ablation Study
Fig. 6 presents the results of using different prompts in our model. It is evident that using the prompt “a vfx with sks” yields superior performance in terms of object shape, color, and other aspects compared to using the category name directly. The reason is that using real object names and defect category names as prompt introduces a significant amount of prior knowledge. However, the prior knowledge from the natural world may differ greatly from the specific objects and defects in the specific dataset. More quantitative results can be found in the supplementary material.
@@ 4.4 Ablation Study
Table 5: Comparison on pixel-level anomaly localization on part of categories in MVTec between with/without the Background Compensation Module. Average is the mean of all categories.
@@
Furthermore, we assess the effectiveness of the Background Compensation Module (BCM), which exhibits more noticeable improvements in the categories of bottle, pill, and toothbrush. The visualization results are depicted in Fig. 4 and the pixel-level localization results are presented in Tab. 5. All the qualitative image generation results and quantitative metrics are improved significantly by using BCM, demonstrating BCM is critical for certain categories.
5 Conclusion
@@
In this paper, we present a novel approach, DualAnoDiff, for generating anomalous image-mask pairs. Our method employs a parallel dual-diffusion to simultaneously generate the anomaly image and the corresponding anomaly part. This ensures a high level of alignment between the generated anomaly image-mask pair and the realism of the anomaly image. Additionally, to address challenging cases, we introduce a background compensation module that efficiently enhances the model’s fitting capability. Extensive experiments demonstrate its superior performance compared to existing anomaly generation methods. The anomaly data generated by our model effectively enhances the performance of downstream anomaly inspection tasks.
Supplementary Material
6 Implementation Details
6.1 Training Details
We train a set of model parameters for each anomaly type. The model requires 5,000 epochs for training, which takes approximately 4.5 hours on an NVIDIA V100 32GB GPU. With BCM, the training step requires only 2,000 epochs. The batch size is set to 4, the learning rate is 0.000005, and the rank of LoRA is 32.
During training, we utilize random flipping for data augmentation. For the “Background Compensation Module”, we have applied it to all categories that involve backgrounds. Among them, categories bottle, pill and toothbrush have witnessed a highly significant improvement. The improvement in the other several categories is rather limited since they can already be generated with a high level of quality.
6.2 Inference Details
During inference without BCM, we only need to input a set of prompts: “a vfx with sks” and “sks”. This process generates a set of anomaly images along with the corresponding anomaly part images. We generate 1000 image pairs with a resolution of 512×512 for each anomaly. Specifically, the num_inference_steps is set to 50, and the guidance_scale is set to 2.5. Notably, it takes 15 seconds to generate each pair of images.
6.3 Mask Generation
We employ U2-Net [30] to segment the anomaly part image and obtain the corresponding mask. Based on our observations, this mask is entirely accurate.
7 More Ablation Studies
We present comprehensive pixel-level and image-level results for downstream anomaly detection in Tables 6 and 7. The term “dual-interrelated diffusion” refers to the utilization of the dual-interrelated model framework, where the type name such as “cable” is employed as a prompt. The notation “+ prompt” indicates the replacement of the type name with “vfx” and “sks”. Additionally, “+BCM” signifies the incorporation of the Background Compensation Module, which is specifically applied to the categories of bottle, grid, hazelnut, pill, and screw. It can be observed that the prompt we designed outperforms the use of category names, with the exception of the toothbrush category. However, the gap between the toothbrush category and the prompt can be effectively bridged by the BCM module.
Category dual-interrelated diffusion $+$prompt $+$prompt $+$BCM
AUC-P AP-P $F_{1}$-P AUC-P AP-P $F_{1}$-P AUC-P AP-P $F_{1}$-P
bottle 96.4 74.2 69.7 98.4 88.8 77.1 99.5 93.4 85.7
cable 95.7 74.1 68.8 97.5 82.6 76.9 97.5 82.6 76.9
capsule 97.8 54.8 54.3 99.5 73.2 67.0 99.5 73.2 67.0
carpet 99.4 86.7 77.9 99.4 89.1 80.2 99.4 89.1 80.2
grid 95.8 36.2 39.8 98.5 57.2 54.9 98.5 57.2 54.9
hazelnut 99.5 94.8 89.9 99.8 96.5 91.5 99.8 97.7 92.8
leather 98.4 79.1 70.1 99.9 88.8 78.8 99.9 88.8 78.8
metal_nut 98.8 94.4 89.1 99.6 98.0 93.0 99.6 98.0 93.0
pill 89.6 38.1 31.2 98.4 86.9 78.2 99.6 95.8 89.2
screw 97.7 48.9 47.9 97.7 55.15 72.8 98.1 57.1 56.1
tile 99.1 91.0 80.8 99.7 97.1 91.0 99.7 97.1 91.0
toothbrush 98.2 65.2 67.1 97.2 62.7 64.0 98.2 68.3 68.6
transistor 94.9 78.2 73.1 98.0 86.7 79.6 98.0 86.7 79.6
wood 98.6 87.3 75.9 99.4 91.6 83.8 99.4 91.6 83.8
zipper 98.4 82.2 72.5 99.6 90.7 82.7 99.6 90.7 82.7
Average 97.22 72.35 67.21 98.8 83.0 78.1 99.1 84.5 78.8
Table 6: Ablaiton Study: comparison on pixel-level anomaly localization on the MVTec dataset by training a U-Net on our model’s generated data using different settings.
Category dual-interrelated diffusion $+$prompt $+$prompt $+$BCM
AUC-P AP-P $F_{1}$-P AUC-I AP-I F1-I AUC-P AP-P $F_{1}$-P
bottle 98.0 99.2 96.4 98.7 98.0 98.9 100 100 100
cable 92.3 94.5 85.1 97.7 98.3 94.2 97.7 98.3 94.2
capsule 81.9 93.5 88.9 97.6 99.2 95.8 97.6 99.2 95.8
carpet 96.7 98.8 95.7 99.8 99.9 99.1 99.8 99.9 99.1
grid 97.2 98.6 95.0 99.5 99.7 97.6 99.5 99.7 97.6
hazelnut 100 100 100 100 100 100 100 100 100
leather 100 100 100 100 100 100 100 100 100
metal_nut 97.7 99.3 97.6 99.7 99.9 99.2 99.7 99.9 99.2
pill 87.1 96.3 91.2 92.0 97.8 93.6 95.8 99.0 95.8
screw 83.5 90.1 84.1 86.6 94.2 86.1 87.8 95.0 87.2
tile 100 100 100 100 100 100 100 100 100
toothbrush 97.9 98.8 94.7 97.6 98.5 93.9 99.5 99.7 97.5
transistor 92.8 92.3 89.4 95.1 93.7 90.1 95.1 93.7 90.1
wood 99.3 99.7 97.6 100 99.9 100 100 99.9 100
zipper 100 100 100 100 100 100 100 100 100
Average 94.9 97.4 94.38 97.6 98.6 96.5 99.8 98.9 99.8
Table 7: Ablaiton Study: comparison on image-level anomaly localization on the MVTec dataset by training a U-Net on our model’s generated data using different settings.
8 More Qualitative Experiments
We conducted a comprehensive comparison between our generated results and those of existing anomaly image generation methods, with the results presented in Fig. 7. It is evident that the diversity of anomalies generated by Crop&Paste [23] is limited. The results from DiffAug [50] exhibit overfitting. The generated outcomes from CDC [27] lack realism, often resulting in distortion, deformation, and other artifacts. SDGAN [26] and Defect-GAN [47] fail to generate masks corresponding to the anomalies, and the authenticity of the generated images is also limited. The masks produced by DFMGAN [6] are not sufficiently aligned, often resulting in the generation of spots or noise. The currently best-performing method, AnomalyDiffusion [14], solely focuses on learning the anomaly part. Consequently, the generated anomaly data fails to integrate smoothly with the original image. And, this sometimes leads to the situation where anomalies manifest against the backdrop of the image. In contrast, our method not only generates highly realistic and diverse anomaly data but also produces highly aligned corresponding masks.
Among all these methods, DFMGAN and AnomalyDiffusion are currently the two best performers, so we conducted a more detailed visualization comparison of our results with these two methods. Additional visualizations are presented in Fig. 8-14. The left side shows two examples from the training data, while the right side displays the generated image pairs.
Figure 7: Comparison on the generation results on MVTec.
Figure 8: Comparison on the type of hazelnut-print. DFMGAN and AnomalyDiffusion struggle to generate realistic anomalies, particularly in the print class, where the anomalous regions consist of strings of letters. In contrast, our method successfully generates both the shape of the letters and the corresponding mask that aligns with their contours.
Figure 9: Comparison on the type of grid-broken. For this specific type of small, structure-related anomaly, the images generated by DFMGAN are of poor quality, and AnomalyDiffusion fails to produce any anomalies. In contrast, our method generates highly realistic and effective anomaly images.
Figure 10: Comparison on the type of toothbrush-defective. The anomalies generated by DFMGAN lack realism, while those produced by AnomalyDiffusion are detached from the main object, Additionally, the generated anomalies, such as holes and bristles of toothbrushes, are mixed. In contrast, although there are some differences in background color, the generated anomalies by our model are fully consistent with real-world scenarios. Furthermore, the background issues do not impact the effectiveness of anomaly detection in downstream tasks.
Figure 11: Comparison on the type of leather-poke. The anomalies generated by AnomalyDiffusion are slightly better than those produced by DFMGAN, however, there is a noticeable color difference in the leather. In contrast, our method achieves good results in both aspects.
Figure 12: Comparison on the type of capsule-scratch. For scratches, a relatively minor type of anomaly, neither DFMGAN nor AnomalyDiffusion can generate effective results. In contrast, our method not only produces realistic anomalies but also demonstrates a good variety.
Figure 13: Comparison on the type of bottle-broken_small. This type of anomaly refers to a small blemish around the edge of a bottle, while broken_large indicates a larger blemish in the same area. The quality of the image generated by DFMGAN is limited, and the mask are not properly aligned. while the abnormal position generated by AnomalyDiffusion sometimes is not correct, and the shape does not belong to the type of broken_small, but more like broken_large. Our method, however, achieves good results in both position and shape.
Figure 14: Comparison on the type of cable-cut_inner_insulation. It is evident that neither DFMGAN nor AnomalyDiffusion can generate realistic anomalies, and the overall quality of the images produced by DFMGAN is subpar. In contrast, our method successfully generates realistic and diverse abnormal data.
9 Quantitative Experiments Setting
9.1 Generated Data
In all comparison methods, 1000 sets of data are generated for each subclass for downstream detection tasks.
9.2 Metrics
This section provides supplementary information on the rationale for using these indicators and their definitions.
For Generation. General image generation tasks typically use Fréchet Inception Distance (FID) [11] to evaluate the difference between the generated data and the real data distribution. However, FID is not reliable in cases of limited anomalous data, as it tends to produce higher scores for overfitted models. Therefore, we utilize the Inception Score (IS) [38] as our evaluation metric. The IS does not require training data and quantifies the quality and diversity of the generated images by calculating the negative exponent of the Kullback-Leibler (KL) divergence between the edge distribution of the generated images and the conditional distribution of the class labels predicted by the Inception model. A higher IS score indicates better quality and diversity in the generated images.
In addition, we use Intra-cluster Pairwise LPIPS Distance (IC-LPIPS) [27] to measure the diversity of the generated data. This method clusters the images into k groups based on the LPIPS distance to k target samples and then computes the average mean LPIPS distances to the corresponding target samples within each cluster. Higher IC-LPIPS scores indicate better diversity.
For Anomaly Inspection. We use the Area Under the Receiver Operating Characteristic (AUROC), Average Precision (AP), and F1-max to measure the performance of the inspection following the general anomaly inspection task.
9.3 Anomaly Inspection Detail
In the downstream task of anomaly detection, we employ a simple U-Net [35] architecture. To mitigate the effects of randomness, we train the model three times and select the best result as the final outcome.